# Periodic Boundary Conditions for Bosonic Path Integral Molecular Dynamics

### v1

@@ II.3 Periodic boundary conditions
Figure 1: Winding configurations for $N=2$ particles in one-dimension with $P=2$ beads (blue for ${\bf r}_{1}^{j}$ and red for ${\bf r}_{2}^{j}$) each and winding cutoff of $\mathcal{W}=2$. In this case, there are only two permutations. The configurations differ by the winding of the spring connecting ${\bf r}_{1}^{P}$ to ${\bf r}_{\sigma(1)}^{1}$ (red solid line), where $\sigma$ is the permutation.
@@ III.1.1 Computing the potential in $\mathcal{O}(\mathcal{W}PN)$ time
Figure 2: An illustration of calculating the distinguishable potential using Equation 15 in the case of a single particle in one-dimension with $P=2$ beads and a winding cutoff of $\mathcal{W}=1$. In this case, there are $\left(2\mathcal{W}+1\right)^{dNP}=9$ different winding configurations surrounded by the dashed line. These configurations are constructed by first summing over windings of individual springs, represented by the “partial” configurations (the contribution from shaded areas are excluded) at the top row and left column, and then multiplying.
@@ III.2.2 Computing the force in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ time
The conditional probability is defined as a sum over all the permutations in which $G[{\sigma}](\ell)=\ell^{\prime}$ and have the winding $\mathbf{w}_{\ell}^{P}$. We show in Section VI of the SI that it can be calculated efficiently by
@@ III.3 Numerical results
Figure 8 shows the convergence with $P$ of both the algorithm with PBC and the MIC approximation for the free Bose gas with $N=64$, $T=0.5\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. If not shown, the statistical error is smaller than the symbol size. We find that MIC converges to the correct result with PBC at about $P=20$. Figure 9 shows the convergence with $P$ for bosons in a sinusoidal trap with $N=32$, $T=1.0\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. We find that the MIC approximation converges to the correct result with PBC at about $P=32$ with the algorithm the rigorously accounts for PBC is converged already for $P=4$. As another example,  Figure 9 compares the results of rigorous PBC treatment with the MIC for the sinusoidal trap. There is a small difference between the PBC algorithm and the MIC approximation in this case, and both converge at about $P=26$.
@@
Figure 11: The probability of all the winding configurations neglected by the MIC approximation during a simulation of $N=32$ bosons in a sinusoidal trap at $T=1.0\;\mathrm{K}$, for values of $P$. As $P$ increases, the probability decays and the MIC approximation becomes better. At $P=14$ MIC and Winding yield practically identical trajectories, which is also when $\expectationvalue{E}/N$ in a MIC simulation converges to the correct result (see Figure 9).

### v2

@@ II.3 Periodic boundary conditions
Figure 1: Winding configurations for $N=2$ particles in one-dimension with $P=2$ beads each (blue for ${\bf r}_{1}^{j}$ and red for ${\bf r}_{2}^{j}$) and winding cutoff of $\mathcal{W}=2$. In this case, there are only two permutations. The configurations differ by the winding of the spring connecting ${\bf r}_{1}^{P}$ to ${\bf r}_{\sigma(1)}^{1}$ (red solid line), where $\sigma$ is the permutation.
@@ III.1.1 Computing the potential in $\mathcal{O}(\mathcal{W}PN)$ time
Figure 2: An illustration of calculating the distinguishable potential using Equation 15 in the case of a single particle in one dimension with $P=2$ beads and a winding cutoff of $\mathcal{W}=1$. In this case, there are $\left(2\mathcal{W}+1\right)^{dNP}=9$ different winding configurations. These configurations are constructed by first summing over windings of individual springs, represented by the “partial” configurations (the contribution from shaded areas are excluded) at the top row and left column, and then multiplying.
@@ III.2.2 Computing the force in $\mathcal{O}(\mathcal{W}(N^{2}+PN))$ time
The conditional probability is defined as a sum over all the configurations in which $G[{\sigma}](\ell)={\ell}^{\prime}$ and have the winding $\mathbf{w}_{\ell}^{P}$. We show in Section VI of the SI that it can be calculated efficiently by
@@ III.3 Numerical results
Figure 8 shows the convergence with $P$ of both the algorithm with PBC and the MIC approximation for the free Bose gas with $N=64$, $T=0.5\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. If not shown, the statistical error is smaller than the symbol size. We find that MIC converges to the correct result with PBC at about $P=14$ while the rigorous treatment of PBC converges at $P=4$. As another example, we compare the results of rigorous PBC treatment with the MIC for the sinusoidal trap. Figure 9 shows the convergence with $P$ for bosons in a sinusoidal trap with $N=32$, $T=1.0\;\mathrm{K}$, $n=0.035\;\text{\AA}^{-3}$. There is a small difference between the PBC algorithm and the MIC approximation in this case, and both converge at about $P=26$.
@@
Figure 11: The probability of all the winding configurations neglected by the MIC approximation during a simulation of $N=32$ bosons in a sinusoidal trap at $T=1.0\;\mathrm{K}$, for different values of $P$. As $P$ increases, the probability decays and the MIC approximation becomes better. At $P=14$ MIC and Winding yield practically identical trajectories, which is also when $\expectationvalue{E}/N$ in a MIC simulation converges to the correct result (see Figure 9).